solve for x.
-6x/5=-7/6
Negative 6x over 5 (as a fraction) = Negative 7 over 6 (as a fraction)

Answer by AnlytcPhil(1806)

solve for x. 
-6x/5=-7/6 
Negative 6x over 5 (as a fraction) = Negative 7 over 6 (as a fraction)

The LCD is 30 because the smallest integer both 5 and 6 will divide
evenly into is 30.  So start with the original equation

        -6x      -7
       ----- = -----
         5        6

Now multiply both sides by 30, written as 30/1:

 30     -6x     30     -7
---- · ----- = ---- · ----
  1      5       1      6

Now on the left cancel the 5 in the bottom into the 30 in the top.
Similarly on the right cancel the 6 in the bottom into the 30 in the top,
Upon cancelling, you get 6 on the left, and 5 on the right:

 6              5 
 30     -6x     30     -7
---- · ----- = ---- · ----
  1      5       1      6
         1              1

Now you have the 6 times the -6x on the left, which gives -36x
Similarly, you have the 5 times the -7 of the right, which gives -35,
so you have:

        -36x = -35   

Now divide both sides by -36. On the left you will have just x, and
on the right side if you divide -35 by -36 you get +35/36

          x = 35/36

Edwin